# Hypothesis testing for community structure in temporal networks using e-values

### v2

@@ 1 Introduction
This paper focuses on testing the null hypothesis of no community structure in a temporally evolving sequence of networks (Holme and Saramäki,, 2012). For the case of a static network, there has been considerable work on this topic (e.g., Lancichinetti et al.,, 2010; Bickel and Sarkar,, 2016; Li and Qi,, 2020; Yuan et al.,, 2022; Yanchenko and Sengupta,, 2024). The key challenged addressed in the aforementioned references is selecting a sensible null model to represent “no community structure,” as well as deriving the null distribution of the test statistic. Indeed, even defining what it means for a network to exhibit community structure is not universally agreed upon (Cazabet and Rossetti,, 2023). For the null model, many methods use the Erdos-Renyi (Erdös and Renyi,, 1959, henceforth ER) model (e.g., Bickel and Sarkar,, 2016) while others use the configuration or Chung-Lu model (Chung and Lu,, 2002; Yanchenko and Sengupta,, 2024). For deriving the null distribution, an asymptotic distribution can be found in some situations (Bickel and Sarkar,, 2016), but when this is not possible, a bootstrap approach is also available (Yanchenko and Sengupta,, 2024).
@@
Despite this extensive work on static networks, to the authors’ knowledge, there are no existing tests for community structure on temporal networks. Users may be tempted to simply aggregate the temporal network into a static network and then use an existing static hypothesis test for community structure. Not only does this remove any temporal variation, but it has also proven to be a poor approach for other network tasks, e.g., Influence Maximization (Erkol et al.,, 2020; Yanchenko et al.,, 2024). Perhaps the closest related work comes from network monitoring and surveillance (Woodall et al.,, 2017; Jeske et al.,, 2018). For example, Wilson et al., (2019) study change-point detection in a temporally evolving network using the degree corrected block model (Karrer and Newman,, 2011, henceforth DCBM). Those authors assume that network snapshots are being generated independently and identically distributed until some unknown time $t^{*}$, at which point the data-generating process changes. The goal is to find this change-point time $t^{*}$. In another work, Wilson et al., (2017) derive a hypothesis test for community structure in multilayer networks, for which temporal networks could be considered as a special case. Their method identifies densely-connected nodes to compute a significance score which is then compared against the configuration model (Newman,, 2006).
@@
When testing on temporal networks, new challenges emerge in addition to inheriting those of the static setting. As in static network testing, carefully defining community structure in a temporal setting is quite difficult (Cazabet and Rossetti,, 2023). Unique to temporal networks on the other hand, sequential realizations induce highly non-trivial dependence, making it difficult to combine results across observations using e.g., p-values.
@@
To overcome these challenges, we develop a community detection hypothesis test for temporal networks using an e-value framework. E-values have recently been gaining popularity in testing settings (e.g., Vovk and Wang,, 2021; Xu and Ramdas,, 2024). The name comes from “expectation” in that their key property is their expectation is less than or equal to 1 when the null hypothesis is true. Thus, large e-values correspond to evidence against the null. Moreover, key properties of e-values address the difficulty of combining results with arbitrary dependence. As many in the network science community may be unfamiliar with e-values, we devote a sizable portion the manuscript to discussing their basic features.
@@
The proposed test first calculates a p-value for (static) community structure on each temporal snapshot before converting this to an e-value. The e-values are then averaged to yield an overall measure of evidence against the null hypothesis of no community structure across networks. Large e-values indicate the likely presence of community structure. By using e-values, we can easily combine information from different snapshots while still maintaining type I error guarantees even with arbitrary dependence in the network process. Additionally, the general nature of our method means it can be applied to a wide-range of data-generating mechanisms and static hypothesis tests. Indeed, we apply the method to two different static hypothesis tests as well as three data-generating mechanisms. Moreover, our experiments on real-data highlight some of the challenges of defining and testing for temporal community structure (Cazabet and Rossetti,, 2023).
@@ 2.2 Challenges in temporal network testing
Hypothesis testing on networks is difficult due to the inherent dependence in the data, and it is made even more difficult when working with temporal networks. Specifically, we need to combine information with unknown and non-trivial dependence. Assume we observe a temporal network $\mathcal{G}=(\mathcal{G}^{(1)},\dots,\mathcal{G}^{(T)})$, consisting of $T$ “snapshots” or realizations of the network-generating process. Each snapshot, $\mathcal{G}^{(t)}$, yields some information about the presence/absence of a community structure in the under-lying data-generating process. We may assume that there are existing methods to extract the evidence of a community structure for each individual snapshot, i.e., using a static hypothesis test for community structure (e.g., Bickel and Sarkar,, 2016; Yanchenko and Sengupta,, 2024). But how can we combine the evidence from each snapshot into an overall measure of evidence of community structure in the temporal network? Indeed, the snapshots likely have a highly non-trivial dependence, greatly complicating their combination. If we use traditional p-values, then in general it is quite challenging to combine this information, especially with unspecified dependence (e.g., Benjamini and Hochberg,, 1995). Therefore, we seek a method that can easily combine information about the underlying data generating structure even with unknown dependence.
@@ 2.3.1 Basics
Before presenting the proposed test, we give some important background on e-values and their basic properties. Recently, e-values have become a popular approach for hypothesis testing (Wasserman et al.,, 2020; Ramdas et al.,, 2020; Dey et al.,, 2024; Xu and Ramdas,, 2024; Grünwald et al.,, 2024). Among other things, they allow for: combination under arbitrary dependence, anytime-valid stopping rules and optional continuation of experiments, but we primarily focus on the combination property. We emphasize that the results in this sub-section are not novel; this discussion is intended to provide the foundation for understanding our proposed method as described in Section 3.
@@ 2.3.2 Combination
We refer the interested reader to Vovk and Wang, (2021); Ramdas and Wang, (2024) for more details on combining e-values. Intuitively, if e-values from the individual snapshots are large, then the arithmetic mean will also be large. In other words, if there is strong evidence for community structure on individual snapshots, then there will be strong evidence for community structure on the temporal network.
@@ 3.1 Proposed test
First, we must compute a p-value on each graph snapshot $\mathcal{G}^{(t)}$ for $t\in\{1,\dots,T\}$. This p-value corresponds to the evidence for/against a community structure on the static component $\mathcal{G}^{(t)}$. We stress that any static hypothesis test for community structure can be used as long as it yields a valid p-value. We primarily use the approach from Bickel and Sarkar, (2016), with details provided in the Supplementary Materials.
@@
Since $\mathcal{G}^{(t)}$ is likely dependent on $\mathcal{G}^{(t^{\prime})}$ for $t^{\prime}\neq t$, the corresponding p-values $P_{t}$ and $P_{t^{\prime}}$ will also be dependent. In general, it is quite difficult to combine dependent p-values, especially if this dependence is unknown. To circumvent this issue, we convert each p-value to an e-value since the latter are trivial to combine. To do this, we use a (p-e) calibrator. A calibrator is a non-negative decreasing function $g:[0,\infty)\to[0,\infty]$ with integral at most one such that $g(x)=0$ for all $x\in(1,\infty)$ and $g(P)$ is an e-value if $P$ is a p-value. In other words, the calibrator transforms a p-value into an e-value; the reverse transformation is also possible, but we will not need this in the present paper. Some examples of calibrators are given in Shafer et al., (2011) and Vovk and Wang, (2021). Specifically, for any $\kappa\in(0,1)$,
@@
and this is a calibrator, called avg. We consider each of these three approaches in the simulation study.
@@
Once the p-values have been converted to e-values, we combine them by taking the arithmetic mean. We showed in Section 2 that not only does this yield a valid e-value, but it can also be shown that this is the best way to combine them with unknown dependence (Vovk and Wang,, 2021). More generally, for any positive weights $w_{1},\dots,w_{T}$ such that $\sum_{t=1}^{T}w_{t}=1$,
@@
After averaging the e-values, we are left with a single e-value which quantifies the evidence for or against the presence of a community structure in the temporal network. Typically in hypothesis testing, the user chooses some threshold $\alpha\in(0,1)$ as their type I error rate, and then rejects if the p-value is smaller than $\alpha$ with common choices $\alpha=0.05$ or $\alpha=0.01$. On the other hand, for e-values, we could reject the null hypothesis when it is “large.” While there is no universally agreed upon threshold, $E>20$ yields a sensible rejection threshold, roughly corresponding to rejecting at $\alpha=0.05$ level (Wang,, 2023). That being said, we do not necessarily endorse the use of 20 as the threshold for decision making. Instead, the raw e-value should be considered as the strength of evidence against the null hypothesis of “no community structure,” and the practitioner can choose the appropriate level, albeit greater than 1.
@@ 3.2 Properties
Theorem 1 follows directly from the basic definition of e-values described in Section 2.3 and allows $\bar{E}_{T}$ to be used for hypothesis testing. The novelty lies not in the mathematical machinery itself, but in the application to testing on temporal networks. Indeed, to the authors’ knowledge, this is the first hypothesis test for community structure in temporal networks.
While not explicitly stated, the proposed test implicitly adopts its null hypothesis of “no community structure” via the static community detection hypothesis test. Many methods use the ER model, e.g., Bickel and Sarkar, (2016); Yuan et al., (2022); Yanchenko and Sengupta, (2024). Here, the probability of an edge between any two nodes is independently $p$, i.e., $A_{ij}\mid p\stackrel{{\scriptstyle\text{iid.}}}{{\sim}}\mathsf{Bernoulli}(p)$. We write ${\bf A}\sim\mathbb{P}_{ER}(p)$ as short-hand to denote the ER data-generating distribution with edge probability $p$. Then $\mathcal{M}_{1}$ as the set of all possible network generating matrices, and $\mathcal{P}_{0}=\{\mathbb{P}_{ER}(p)\mid p\in(0,1)\}$. Tautologically, this model does not encode any community structure. Furthermore, certain data-generating models implicitly encode the ER model as a null model (e.g., Sasahara et al.,, 2021).
Practitioners may be wary of this choice of null model. In particular, the model may be so simple that virtually all real-world networks diverge from it, whether they possess a community structure or not. In such cases, the configuration model may be a more appropriate null model (Lancichinetti et al.,, 2010; Palowitch et al.,, 2017). For ease of exposition, we focus on the ER null model and Bickel and Sarkar, (2016) hypothesis testing framework, but we stress that the method can easily be generalized to other null models and/or hypothesis tests. Indeed, any community structure hypothesis test for static networks can be adapted to this framework. We explore such generalizations in the real-data analysis.
As for the alternative hypothesis, this is generally not explicitly encoded in the hypothesis test. Instead, the definition of “community structure” will be determined by the user depending on their data-generating model. A sensible choice would be the stochastic block model (SBM) (Holland et al.,, 1983) but it can be more general. As our preferred hypothesis test from Bickel and Sarkar, (2016) does not encode a specific alternative hypothesis, the proposed test does not make any modeling assumptions of the network-generating process. This means that the test can be used for a wide range of models and settings, e.g., the community structure changes over time as nodes change groups, new communities form and/or disappear, new nodes are added to the network, etc. Similarly, we did not need to specify the dependence between successive observations $\mathcal{G}^{(t)}$ and $\mathcal{G}^{(t+1)}$ because the averaging of e-values is still a valid e-value, regardless of their dependence.
@@ 4.1 Set-up
In this section, we perform numerical simulations to study the properties of the proposed hypothesis test, primarily demonstrating its ability to combine results across dependent network snapshots. Thus, we assume that the number of observations, $T$, was fixed before the experiment began.
As the test is not tied to a particular data-generating model, we consider three different network-generating models. For each model, we vary parameter settings which control the strength of the community structure or the number of nodes, and report the median e-value over 100 Monte Carlo (MC) iterations. To clarify, the individual e-values are calculated as the average e-value from each snapshot, while the median is taken over the MC replicates. We compare the proposed test using five different calibrators: max, avg and $\kappa=0.25,0.50,0.75$. Recall that max is not a proper calibrator but we may refer to it as one for ease of exposition. We prefer the median as opposed to the mean as an infinite e-value from just a single MC iteration can obscure the overall trends. For the variability results of the e-values, please see the Supplemental Materials. We stress that he goal of these simulations is not to identify the specific communities, but rather to test for the presence/absence of a community structure. Finally, code to implement the proposed test and reproduce the simulation studies is available on the author’s GitHub: https://github.com/eyanchenko/tempComDet#.
@@ 4.2.1 Model
First, we propose a novel model to generate correlated observations with a stochastic block model (Holland et al.,, 1983). The stochastic block model (SBM) is a popular model to generate networks with a community structure in static networks. We present the basic ideas before generalizing to temporal networks. Given some network with $n$ nodes and $K$ communities, let $\boldsymbol{c}\in\{1,\dots,K\}^{n}$ be such that $c_{i}=k$ if node $i$ is in community $k$. Here, $\boldsymbol{c}$ represents the latent community structure. Given $\boldsymbol{c}$, the probability of an edge between nodes $i$ and $j$, $P_{ij}$, is
$P_{ij}=B_{c_{i},c_{j}}$
where ${\bf B}\in(0,1)^{K\times K}$ is the block-probability matrix such that $B_{k,k^{\prime}}$ is the probability of an edge between nodes in community $k$ and $k^{\prime}$ for $k,k^{\prime}\in\{1,\dots,K\}$. If $B_{k,k^{\prime}}=b$ for all $k,k^{\prime}\in\{1,\dots,K\}$, then this model reduces to the ER (null) model. On the other hand, if the diagonal entries of $\bf B$} are larger than those on the off-diagonal, then the SBM generates networks with assortative community structure (alternative hypothesis).
To extend this to the temporal setting, we are inspired by the graph matching literature (see, e.g. Lyzinski et al.,, 2014) to generate a (potentially correlated) sequence of network realizations, or snapshots. For the first time step, $t=1$, we simply generate a static SBM network, i.e.,
@@
$A_{ij}^{(t)}|A_{ij}^{(t-1)},\boldsymbol{c}\stackrel{{\scriptstyle ind.}}{{\sim}}\begin{cases}\mathsf{Bernoulli}\{B_{c_{i},c_{j}}+\rho(1-B_{c_{i},c_{j}})\}&\text{if }A_{ij}^{(t-1)}=1\\ \mathsf{Bernoulli}\{B_{c_{i},c_{j}}(1-\rho)\}&\text{if }A_{ij}^{(t-1)}=0\end{cases}$ (3)
This formulation ensures that
$\mathsf{Cor}(A_{ij}^{(t)},A_{ij}^{(t-1)})=\rho.$
Please see the Supplemental Materials for proof. As special cases, if $\rho=0$, then the snapshots are independent, and if $\rho=1$, then they are identical. This resembles a first-order Markov process as the next snapshot only depends on the current one, and $\rho$ controls the amount of correlation between the successive realizations. Additionally, this model assumes that the latent community labels, $\boldsymbol{c}$, do not vary with time, but we relax this assumption in the subsequent model.

### v3

@@ 1 Introduction
This paper focuses on testing the null hypothesis of no community structure in a temporally evolving sequence of networks (Holme and Saramäki, 2012). For the case of a static network, there has been considerable work on this topic (e.g., Lancichinetti et al., 2010; Bickel and Sarkar, 2016; Li and Qi, 2020; Yuan et al., 2022; Yanchenko and Sengupta, 2024). The key challenged addressed in the aforementioned references is selecting a sensible null model to represent “no community structure,” as well as deriving the null distribution of the test statistic. Indeed, even defining what it means for a network to exhibit community structure is not universally agreed upon (Cazabet and Rossetti, 2023). For the null model, many methods use the Erdos-Renyi (Erdös and Renyi, 1959, henceforth ER) model (e.g., Bickel and Sarkar, 2016) while others use the configuration or Chung-Lu model (Chung and Lu, 2002; Yanchenko and Sengupta, 2024). For deriving the null distribution, an asymptotic distribution can be found in some situations (Bickel and Sarkar, 2016), but when this is not possible, a bootstrap approach is also available (Yanchenko and Sengupta, 2024).
@@
Despite this extensive work on static networks, to the authors’ knowledge, there are no existing tests for community structure on temporal networks. Users may be tempted to simply aggregate the temporal network into a static network and then use an existing static hypothesis test for community structure. Not only does this remove any temporal variation, but it has also proven to be a poor approach for other network tasks, e.g., Influence Maximization (Erkol et al., 2020; Yanchenko et al., 2024). Perhaps the closest related work comes from network monitoring and surveillance (Woodall et al., 2017; Jeske et al., 2018). For example, Wilson et al. (2019) study change-point detection in a temporally evolving network using the degree corrected block model (Karrer and Newman, 2011, henceforth DCBM). Those authors assume that network snapshots are being generated independently and identically distributed until some unknown time $t^{*}$, at which point the data-generating process changes. The goal is to find this change-point time $t^{*}$. In another work, Wilson et al. (2017) derive a hypothesis test for community structure in multilayer networks, for which temporal networks could be considered as a special case. Their method identifies densely-connected nodes to compute a significance score which is then compared against the configuration model (Newman, 2006).
@@
When testing on temporal networks, new challenges emerge in addition to inheriting those of the static setting. As in static network testing, carefully defining community structure in a temporal setting is quite difficult (Cazabet and Rossetti, 2023). Unique to temporal networks on the other hand, sequential realizations induce highly non-trivial dependence, making it difficult to combine results across observations using e.g., p-values.
@@
To overcome these challenges, we develop a community detection hypothesis test for temporal networks using an e-value framework. E-values have recently been gaining popularity in testing settings (e.g., Vovk and Wang, 2021; Xu and Ramdas, 2024). The name comes from “expectation” in that their key property is their expectation is less than or equal to 1 when the null hypothesis is true. Thus, large e-values correspond to evidence against the null. Moreover, key properties of e-values address the difficulty of combining results with arbitrary dependence. As many in the network science community may be unfamiliar with e-values, we devote a sizable portion the manuscript to discussing their basic features.
@@
The proposed test first calculates a p-value for (static) community structure on each temporal snapshot before converting this to an e-value. The e-values are then averaged to yield an overall measure of evidence against the null hypothesis of no community structure across networks. Large e-values indicate the likely presence of community structure. By using e-values, we can easily combine information from different snapshots while still maintaining type I error guarantees even with arbitrary dependence in the network process. Additionally, the general nature of our method means it can be applied to a wide-range of data-generating mechanisms and static hypothesis tests. Indeed, we apply the method to two different static hypothesis tests as well as three data-generating mechanisms. Moreover, our experiments on real-data highlight some of the challenges of defining and testing for temporal community structure (Cazabet and Rossetti, 2023).
@@ 2.2 Challenges in temporal network testing
Hypothesis testing on networks is difficult due to the inherent dependence in the data, and it is made even more difficult when working with temporal networks. Specifically, we need to combine information with unknown and non-trivial dependence. Assume we observe a temporal network $\mathcal{G}=(\mathcal{G}^{(1)},\dots,\mathcal{G}^{(T)})$, consisting of $T$ “snapshots” or realizations of the network-generating process. Each snapshot, $\mathcal{G}^{(t)}$, yields some information about the presence/absence of a community structure in the under-lying data-generating process. We may assume that there are existing methods to extract the evidence of a community structure for each individual snapshot, i.e., using a static hypothesis test for community structure (e.g., Bickel and Sarkar, 2016; Yanchenko and Sengupta, 2024). But how can we combine the evidence from each snapshot into an overall measure of evidence of community structure in the temporal network? Indeed, the snapshots likely have a highly non-trivial dependence, greatly complicating their combination. If we use traditional p-values, then in general it is quite challenging to combine this information, especially with unspecified dependence (e.g., Benjamini and Hochberg, 1995). Therefore, we seek a method that can easily combine information about the underlying data generating structure even with unknown dependence.
@@ 2.3.1 Basics
Before presenting the proposed test, we give some important background on e-values and their basic properties. Recently, e-values have become a popular approach for hypothesis testing (Wasserman et al., 2020; Ramdas et al., 2020; Dey et al., 2024; Xu and Ramdas, 2024; Grünwald et al., 2024). Among other things, they allow for: combination under arbitrary dependence, anytime-valid stopping rules and optional continuation of experiments, but we primarily focus on the combination property. We emphasize that the results in this sub-section are not novel; this discussion is intended to provide the foundation for understanding our proposed method as described in Section 3.
@@ 2.3.2 Combination
We refer the interested reader to Vovk and Wang (2021); Ramdas and Wang (2024) for more details on combining e-values. Intuitively, if e-values from the individual snapshots are large, then the arithmetic mean will also be large. In other words, if there is strong evidence for community structure on individual snapshots, then there will be strong evidence for community structure on the temporal network.
@@ 3.1 Proposed test
First, we must compute a p-value on each graph snapshot $\mathcal{G}^{(t)}$ for $t\in\{1,\dots,T\}$. This p-value corresponds to the evidence for/against a community structure on the static component $\mathcal{G}^{(t)}$. We stress that any static hypothesis test for community structure can be used as long as it yields a valid p-value. We primarily use the approach from Bickel and Sarkar (2016), with details provided in the Supplementary Materials.
@@
Since $\mathcal{G}^{(t)}$ is likely dependent on $\mathcal{G}^{(t^{\prime})}$ for $t^{\prime}\neq t$, the corresponding p-values $P_{t}$ and $P_{t^{\prime}}$ will also be dependent. In general, it is quite difficult to combine dependent p-values, especially if this dependence is unknown. To circumvent this issue, we convert each p-value to an e-value since the latter are trivial to combine. To do this, we use a (p-e) calibrator. A calibrator is a non-negative decreasing function $g:[0,\infty)\to[0,\infty]$ with integral at most one such that $g(x)=0$ for all $x\in(1,\infty)$ and $g(P)$ is an e-value if $P$ is a p-value. In other words, the calibrator transforms a p-value into an e-value; the reverse transformation is also possible, but we will not need this in the present paper. Some examples of calibrators are given in Shafer et al. (2011) and Vovk and Wang (2021). Specifically, for any $\kappa\in(0,1)$,
@@
the calibrator we call avg. We consider each of these three approaches in our simulations.
@@
Once the p-values have been converted to e-values, we combine them by taking the arithmetic mean. We showed in Section 2 that not only does this yield a valid e-value, but it can also be shown that this is the best way to combine them with unknown dependence (Vovk and Wang, 2021). More generally, for any positive weights $w_{1},\dots,w_{T}$ such that $\sum_{t=1}^{T}w_{t}=1$,
@@
After averaging the e-values, we are left with a single e-value which quantifies the evidence for or against the presence of a community structure in the temporal network. Typically in hypothesis testing, the user chooses some threshold $\alpha\in(0,1)$ as their type I error rate, and then rejects if the p-value is smaller than $\alpha$ with common choices $\alpha=0.05$ or $\alpha=0.01$. On the other hand, for e-values, we could reject the null hypothesis when it is “large.” While there is no universally agreed upon threshold, $E>20$ yields a sensible rejection threshold, yielding a type I error of no more than $\alpha=0.05$ (Wang, 2023). That being said, we do not necessarily endorse the use of 20 as the threshold for decision making. Instead, the raw e-value should be considered as the strength of evidence against the null hypothesis of “no community structure,” and the practitioner can choose the appropriate level, albeit greater than 1.
@@ 3.2 Properties
Theorem 1 follows directly from the basic definition of e-values described in Section 2.3 and allows $\bar{E}_{T}$ to be used for hypothesis testing. The novelty here lies in the application to testing on temporal networks. Indeed, to the authors’ knowledge, this is the first hypothesis test for community structure in temporal networks.
While not explicitly stated, the proposed test implicitly adopts its null hypothesis of “no community structure” via the static community detection hypothesis test. Many methods use the ER model, e.g., Bickel and Sarkar (2016); Yuan et al. (2022); Yanchenko and Sengupta (2024). Here, the probability of an edge between any two nodes is independently $p$, i.e., $A_{ij}\mid p\stackrel{{\scriptstyle\text{iid.}}}{{\sim}}\mathsf{Bernoulli}(p)$. We write ${\bf A}\sim\mathbb{P}_{ER}(p)$ as short-hand to denote the ER data-generating distribution with edge probability $p$. Then $\mathcal{M}_{1}$ as the set of all possible network generating matrices, and $\mathcal{P}_{0}=\{\mathbb{P}_{ER}(p)\mid p\in(0,1)\}$. Tautologically, this model does not encode any community structure. Furthermore, certain data-generating models implicitly encode the ER model as a null model (e.g., Sasahara et al., 2021).
Practitioners may be wary of this choice of null model. In particular, the model may be so simple that virtually all real-world networks diverge from it, whether they possess a community structure or not. In such cases, the configuration model may be a more appropriate null model (Lancichinetti et al., 2010; Palowitch et al., 2017). For ease of exposition, we focus on the ER null model and Bickel and Sarkar (2016) hypothesis testing framework, but we stress that the method can easily be generalized to other null models and/or hypothesis tests. Indeed, any community structure hypothesis test for static networks can be adapted to this framework. We explore such generalizations in the real-data analysis.
As for the alternative hypothesis, this is generally not explicitly encoded in the hypothesis test. Instead, the definition of “community structure” will be determined by the user depending on their data-generating model. A sensible choice would be the stochastic block model (SBM) (Holland et al., 1983) but it can be more general. As our preferred hypothesis test from Bickel and Sarkar (2016) does not encode a specific alternative hypothesis, the proposed test does not make any modeling assumptions of the network-generating process. This means that the test can be used for a wide range of models and settings, e.g., the community structure changes over time as nodes change groups, new communities form and/or disappear, new nodes are added to the network, etc. Similarly, we did not need to specify the dependence between successive observations $\mathcal{G}^{(t)}$ and $\mathcal{G}^{(t+1)}$ because the averaging of e-values is still a valid e-value, regardless of their dependence.
@@ 4.1 Set-up
Here we perform numerical simulations to study the properties of the proposed hypothesis test, demonstrating its ability to combine results across dependent network snapshots. Thus, we assume that the number of observations, $T$, was fixed before the experiment began.
As the test is not tied to a particular data-generating model, we consider three different network-generating models. Since we use the static test from Bickel and Sarkar (2016), for each model, the hypotheses being tested are
[…]
We vary parameter settings which control the strength of the community structure or the number of nodes, and report the median e-value over 100 Monte Carlo (MC) iterations. To clarify, the individual e-values are calculated as the average e-value from each snapshot, while the median is taken over the MC replicates. We compare the proposed test using five different calibrators: max, avg and $\kappa=0.25,0.50,0.75$. Recall that max is not a proper calibrator but we may refer to it as one for ease of exposition. We prefer the median as opposed to the mean as an infinite e-value from just a single MC iteration can obscure the overall trends. While we are primarily interested in reporting the e-values of the test directly, we also report the type I error and power of the test for a rejection threshold of $\bar{E}_{T}>20$ in the Supplemental Materials, in addition to variability results. We stress that he goal of these simulations is not to identify the specific communities, but rather to test for the presence/absence of a community structure. Finally, code to implement the proposed test and reproduce the simulation studies is available from the first author’s GitHub page: https://github.com/eyanchenko/tempComDet#.
@@ 4.2.1 Model
First, we propose a novel model to generate correlated observations with a stochastic block model (Holland et al., 1983). The stochastic block model (SBM) is a popular model to generate networks with a community structure in static networks. We present the basic ideas before generalizing to temporal networks. Given some network with $n$ nodes and $K$ communities, let $\boldsymbol{c}\in\{1,\dots,K\}^{n}$ be such that $c_{i}=k$ if node $i$ is in community $k$. Here, $\boldsymbol{c}$ represents the latent community structure. Given $\boldsymbol{c}$, the probability of an edge between nodes $i$ and $j$, $P_{ij}$, is $P_{ij}=B_{c_{i},c_{j}}$, where ${\bf B}\in(0,1)^{K\times K}$ is the block-probability matrix such that $B_{k,k^{\prime}}$ is the probability of an edge between nodes in community $k$ and $k^{\prime}$ for $k,k^{\prime}\in\{1,\dots,K\}$. If $B_{k,k^{\prime}}=b$ for all $k,k^{\prime}\in\{1,\dots,K\}$, then this model reduces to the ER (null) model. On the other hand, if the diagonal entries of $\bf B$} are larger than those on the off-diagonal, then the SBM generates networks with assortative community structure (alternative hypothesis).
To extend this to the temporal setting, we are inspired by the graph matching literature (see, e.g. Lyzinski et al., 2014) to generate a (potentially correlated) sequence of network realizations, or snapshots. At time step, $t=1$, we generate a static SBM network, i.e.,
@@
$A_{ij}^{(t)}|A_{ij}^{(t-1)},\boldsymbol{c},{\bf B}\stackrel{{\scriptstyle ind.}}{{\sim}}\begin{cases}\mathsf{Bernoulli}\{B_{c_{i},c_{j}}+\rho(1-B_{c_{i},c_{j}})\}&\text{if }A_{ij}^{(t-1)}=1\\ \mathsf{Bernoulli}\{B_{c_{i},c_{j}}(1-\rho)\}&\text{if }A_{ij}^{(t-1)}=0\end{cases}$ (3)
This formulation ensures that $\mathsf{Cor}(A_{ij}^{(t)},A_{ij}^{(t-1)})=\rho$, which we prove in the Supplemental Materials. As special cases, if $\rho=0$, then the snapshots are independent, and if $\rho=1$, then they are identical. This resembles a first-order Markov process as the next snapshot only depends on the current one, and $\rho$ controls the amount of correlation between the successive realizations. Additionally, this model assumes that the latent community labels, $\boldsymbol{c}$, do not vary with time, but we relax this assumption in the subsequent model.